OpenStudy (christos):

Whats the integral of ln(x^2) ?

OpenStudy (zzr0ck3r):

ln(x^2) = 2ln(x)

OpenStudy (zzr0ck3r):

what is the integral of ln(x)?

OpenStudy (christos):

1/x ?

OpenStudy (zzr0ck3r):

no that's the derivative

OpenStudy (christos):

wait

OpenStudy (zzr0ck3r):

hint: let u = ln(x) and let dv = 1 dx

OpenStudy (zzr0ck3r):

try not to just give answers? we can go to wolfram for that

OpenStudy (primeralph):

@zzr0ck3r still not the final answer; but I'll take it off

OpenStudy (christos):

xln(x) ? I still refuse to go by the answer :D

OpenStudy (primeralph):

@Christos xlnx is not the answer. Use integration by parts

OpenStudy (christos):

xln(x)-1

OpenStudy (zzr0ck3r):

if you are just guessing you can go to wolframalpha and get the answer

OpenStudy (christos):

I am not guessing I though x has a power of 1 so -1 at the end and outside of ln whatever is inside

OpenStudy (christos):

I saw it somewhere long time ago but im not sure if its correct

OpenStudy (zzr0ck3r):

u need to do integration by parts

OpenStudy (zzr0ck3r):

is that what you are doing in class?

OpenStudy (christos):

What is this, can you give me an example?

OpenStudy (christos):

I am learning alone.

OpenStudy (zzr0ck3r):

do you know what the product rule is with derivatives?

OpenStudy (christos):

I do

OpenStudy (zzr0ck3r):

it is like the inverse of that, it is how we undo that rule.

OpenStudy (zzr0ck3r):

I would google it and read on it, it takes practice

OpenStudy (christos):

the reverse?

OpenStudy (zzr0ck3r):

sort of

OpenStudy (zzr0ck3r):

do you know u substitution for integration?

OpenStudy (christos):

yes

OpenStudy (zzr0ck3r):

ok that "undoes" the chain rule, so now you need to learn by parts. there are both crucial to learning integration http://tutorial.math.lamar.edu/Classes/CalcII/IntegrationByParts.aspx

OpenStudy (christos):

Look at this please: That's specifically what am trying to solve. http://screencast.com/t/0N48E4lhqM9

OpenStudy (zzr0ck3r):

e^(2*lnx) = e^ln(x^2) = x^2 so you need the integral of x^2 not ln(x^2)

OpenStudy (christos):

I solved it now :D I know how to find this thing very easy (x^3)/3 without any actual formula

OpenStudy (zzr0ck3r):

yeah:)

OpenStudy (christos):

thanks

OpenStudy (zzr0ck3r):

np